The supplementary of an angle measures 25 degrees more than it's compliment.find the measure of the angle
Answer by Alan3354(69443)

The supplementary of an angle measures 25 degrees more than it's compliment.find the measure of the angle
------------
It's the supplement of an angle, not the supplementary.
It's complement, not compliment.
it's = it is.
--------------
a = the angle
180 - a = the supplement
90 - a = the complement
180 - a = 90 - a + 25
180 = 115
No solution.
--------------
The supplement of an angle is always 90 deg more than the complement.